OpenStudy (anonymous):

Write an equation of the line passes through the points:(4,2),(-2,-4)

OpenStudy (anonymous):

(2+4)/(4+2)= slope or m -> 6/6 or 1

OpenStudy (anonymous):

Ohh okay so basically im just finding the slope?

OpenStudy (anonymous):

2=1(4)+b b=-2 using this info y=1x-2

OpenStudy (anonymous):

okay so first i find the slope then just plug that into the equation?

OpenStudy (anonymous):

you find the slope first. Then you plug in the given x and y values into the y=mx+B equation in order to find b

OpenStudy (anonymous):

okay xD

OpenStudy (anonymous):

m=slope b=y intercept

OpenStudy (anonymous):

k i think i get how to do it them now, thx

OpenStudy (anonymous):

you know how to find the slope between 2 points right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok, and it doesn't matter what x and y value you plug in as long as their in the same set like how I plugged in x=4 and y=2 (4,2)

OpenStudy (anonymous):

so you could plug (-2,-4) in also?